OpenStudy (karatechopper):

Tim's bank contains quarters, dimes, and nickles. He has three more dimes than quarters and 6 fewer nickles then quarters. If tim has 63 coins write and solve an equation to find how many quarters tim has

OpenStudy (anonymous):

d+q+n=63 3q+q+(q-6)=63 solve for q

OpenStudy (karatechopper):

how can that be 3q we r talking dimes and quarters should tht not be a 3d?

OpenStudy (anonymous):

sorry d is dime q is quarters and n is nickles (3+q)+q+(q-6)=63 q+q+q=66 3q=66 q=33

OpenStudy (karatechopper):

what about the 3 and 6?

OpenStudy (karatechopper):

wouldn't it be 9 + 3q?

OpenStudy (anonymous):

which 3 and 6

OpenStudy (karatechopper):

3 for more dimes and 6 for nickles

OpenStudy (anonymous):

no it is -6 so it is 3-6 which is -3

OpenStudy (karatechopper):

so -3q=63?

OpenStudy (anonymous):

no so you have 3+q+q+q-6=63 3q-3=63 3q=63+3 3q=66 q=33

OpenStudy (karatechopper):

it actually 22 not 33

OpenStudy (anonymous):

yeah sorry

OpenStudy (karatechopper):

ok medal and fan 4 u then!

OpenStudy (anonymous):

Solve the following equations for d,n and q:\[\{d=q+3,n=q-6,q+n+d=63\}\]\[\{q=22,n=16,d=25\} \]